Consider a pointer pointing to an array A. int A[
Answer #1

a) 14 (initial value of pointer-- *P:12-- + 2)
b) 34 (*(P+2) makes the pointer shift to initial + 2 position, i.e. pointer moves from first to third posiotion, value here is 34)
c) memory address of pointers current position + 4 bytes of int
d) memory address of A[4]=56-3=53 (53 temporary value in memory)
e) memory Address of A[0] initial pointer + 4*3 = 12 bytes
f) 7
g) memory address of pointer P - 4*10 = 4= bytes
h) 23


In the given program, the method GetSinCos Doesn't return any values so the values of Sin(dx) and Cos(dx) remains in the variable dSin and dCos declared within the method GetSinCos and is not accessible in main method.To get the correct values, you can use the cout statements within the GetSinCos method instead of using it in main method.


answered by: ANURANJAN SARSAM